# Subtree Mode and Applications

### v1

@@ 9 Experimental Evaluation
From each dataset, we mined uniform patterns and show the $3$ patterns with the largest difference among their frequencies in the two strings of the dataset. Table 3 shows the patterns, their frequency in each string of their dataset, the $\varepsilon$ value used, and references supporting that indeed each pattern is prevalent in the subpopulation in which it has the largest frequency in our dataset. These patterns effectively reveal behavioral differences between different subpopulations. In Movielens, they show which movie genres are preferred mostly by men (first and third pattern) or by women (second pattern). These movie genres are indeed preferred by the respective subpopulations [WLS17, KK05, Inf+21]. In Book-Crossing, the patterns show which book genres are preferred mostly by elderly (first pattern) or by teenagers (second and third pattern) and indeed this agrees with the literature [MK ̵14, CWM25, Inf+21, Dub+23]. Last, in Alibaba, the patterns show that customers with high (respectively, low) purchase power browse products in the high or premium price tier (respectively, in the low price tier), and this is again well-supported by the literature [HND10, NFT19, Lu20, MHV17, BW10].
@@
We used the SARS-CoV-2 [Nat24] (SARS) database as $\mathcal{S}$. SARS contains $2,000$ strings, each representing a different genetic variation of SARS-CoV-2. The total length of strings in SARS is $59,515,733$ and the alphabet size of all these strings is $4$. We randomly selected $50$ strings from SARS as queries $Q$ and removed them from the database. Additionally, we retrieved $50$ Influenza A genomes [Nat25] (INFL) and sampled $50$ substrings from the human genome [Nat22] (HUM). Each retrieved string from INFL and each sampled substring of HUM was used as a query on SARS. Since the genome size of INFL is smaller than that of the SARS virus ($\approx 13$kb and $\approx 30$kb, respectively), we used the entire Influenza A genomes as queries. On the other hand, the genome of HUM is much larger ($\approx 3$Gb), so we sampled HUM substrings that were of very similar length to the SARS genome. In particular, the average length of the SARS, INFL, and HUM queries is 29,777, 13,606, and 29,758, respectively, and all queries have roughly the same number of distinct $q$-grams for all tested $q$ values. We used $\varepsilon\in\{0,1,2\}$ and measured the percentage of $\varepsilon$-consistent $q$-grams for different length of the query $q\in[4,7]$.

### v2

@@ 9 Experimental Evaluation
From each dataset, we mined uniform patterns and show the $3$ patterns with the largest difference among their frequencies in the two strings of the dataset. Table 3 shows the patterns, their frequency in each string of their dataset, the $\varepsilon$ value used, and references supporting that indeed each pattern is prevalent in the subpopulation in which it has the largest frequency in our dataset. These patterns effectively reveal behavioral differences between different subpopulations. In Movielens, they show which movie genres are preferred mostly by men (first and third pattern) or by women (second pattern). These movie genres are indeed preferred by the respective subpopulations [wuhr2017tears, krcmar2005uses, infortuna2021inner]. In Book-Crossing, the patterns show which book genres are preferred mostly by elderly (first pattern) or by teenagers (second and third pattern) and indeed this agrees with the literature [awriterofhistory_2014, currie2025reading, infortuna2021inner, dubourg2023exploratory]. Last, in Alibaba, the patterns show that customers with high (respectively, low) purchase power browse products in the high or premium price tier (respectively, in the low price tier), and this is again well-supported by the literature [han2010signaling, ngwe2019impact, lu2020status, mundel2017exploratory, berger2010subtle].
@@
We used the SARS-CoV-2 [ncbi_sars_cov_2_2024] (SARS) database as $\mathcal{S}$. SARS contains $2,000$ strings, each representing a different genetic variation of SARS-CoV-2. The total length of strings in SARS is $59,515,733$ and the alphabet size of all these strings is $4$. We randomly selected $50$ strings from SARS as queries $Q$ and removed them from the database. Additionally, we retrieved $50$ Influenza A genomes [ncbi_influenza_a_2025] (INFL) and sampled $50$ substrings from the human genome [ncbi_human_genome_2022] (HUM). Each retrieved string from INFL and each sampled substring of HUM was used as a query on SARS. Since the genome size of INFL is smaller than that of the SARS virus ($\approx 13$kb and $\approx 30$kb, respectively), we used the entire Influenza A genomes as queries. On the other hand, the genome of HUM is much larger ($\approx 3$Gb), so we sampled HUM substrings that were of very similar length to the SARS genome. In particular, the average length of the SARS, INFL, and HUM queries is 29,777, 13,606, and 29,758, respectively, and all queries have roughly the same number of distinct $q$-grams for all tested $q$ values. We used $\varepsilon\in\{0,1,2\}$ and measured the percentage of $\varepsilon$-consistent $q$-grams for different length of the query $q\in[4,7]$.
@@
We compared our approach to a state-of-the-art method for viral sequence classification [vir3], which uses vectors of $q$-gram frequencies as features to train random forests. The method, referred to as RF, was implemented in Python using scikit-learn and configured and trained as in [vir3]. As RF requires also negative samples (i.e., also INFL and HUM strings in its training set), we used two training sets: (1) $T_{1}$, comprised of $3,950$ strings; the $1,950$ SARS strings we used in the experiment of Fig. 11, $1000$ randomly selected INFL strings of average length $13208$ from [ncbi_influenza_a_2025], and $1000$ randomly selected HUM strings of average length $29758$ from the human genome [ncbi_human_genome_2022]; and (2) $T_{2}$ constructed from $T_{1}$ by removing $950$, $500$, and $500$ randomly selected SARS, INFL, and HUM strings, respectively. SARS strings had a label True and the others a label False. Table 4(a) shows the time for training RF on $T_{1}$ or $T_{2}$ and the time for our approach, for varying $q$. As expected, RF takes more time on $T_{1}$ and when $q$ is larger, while our approach is insensitive to $q$ and significantly faster. In fact, when we increased the training set size and $q$, our approach was up to $30$ times faster. This can be seen in Table 4(b); the training set $T_{3}$ for RF was comprised of $20000$ randomly selected strings from SARS with average length $29773$, which were also used by our approach, and also $10000$ randomly selected strings from the human genome [ncbi_human_genome_2022] with average length $29758$. Furthermore, RF used up to $3.9$ times more memory. In all experiments, we used the queries of the experiment of Fig. 11 for both RF and our approach, $\varepsilon=0$, and classified every query as SARS when $99\%$ of its $q$-grams are $\varepsilon$-consistent and as not-SARS otherwise. Both RF and our approach had no misclassification errors.
$q$ RF RF Our approach
$T_{1}$ $T_{2}$ SARS
44 87.15 46.22 16.57
55 102.06 52.97 16.57
66 119.41 60.82 16.57
77 154.26 77.67 16.57
$q$ RF Our approach
$T_{3}$ $20,000$ SARS
66 501.24 67.74
77 657.33 67.74
88 1022.48 67.74
99 2128.33 67.74
Table 4: (a) Runtime (seconds) for RF trained on $T_{1}$ or $T_{2}$ and for SCM using the SARS database. (b) Runtime (seconds) for RF trained on $T_{3}$ and for SCM using $20,000$ SARS strings.
(a) SARS
(b) SARS
(c) SARS
(d) SARS
(e) SARS
(f) SARS
Figure 12: Runtime and memory for SARS vs. (a-b) $k$, (c-d) NN for $k=8$, and (e-f) $\Delta$, for $k=8$ and $N=7,439,466$.
$k$-DR.
We applied the modified SCM algorithm (see Corollary 2) on the SARS database. Recall from Section 8 that there are no practical methods for $k$-DR (or $k$-SM). Also, it is unclear if and how these problems could be addressed based on BA2. Thus, we compared our algorithm to: (1) a modified version of BA1, which derives the sorted list $(c_{1},f_{1}),\ldots,(c_{k},f_{k})$ per node $v$ based on array $A_{v}$ (see Corollary 2 and Proposition 1), and it takes $\mathcal{O}(N\Delta)$ time and space, and (2) a modified version of BA3 following Corollary 2 which takes $\mathcal{O}(N\log\Delta+kN)$ time and $\mathcal{O}(kN)$ space. Fig. 12(a) shows that, for all $k$ values, our algorithm is faster by at least one order of magnitude than the modified BA3 and even faster than the modified BA1. It also uses at least $6.4\%$ and $75.7\%$ less memory than the modified BA3 and BA1, respectively (see Fig. 12(b)). For varying $N$ (respectively, $\Delta$), our algorithm is $11$ (respectively, $10$) times faster on average than the fastest baseline (see Figs. 12(c) and 12(e)), and it uses $15.3\%$ (respectively, $13.8\%$) less memory on average (see Figs. 12(d) and 12(f)). All results are in line with the complexities of the algorithms and show that our algorithm is the most time- and space-efficient.
Phylogenetic Tree Annotation.
We show that our SCM algorithm helps detecting candidate genotype-phenotype links. We used the phylogenetic tree of [10193, Fig. S1] which is constructed on Treponema pallidium (a syphilis-causing bacterium) data from the study of [Arora2017OriginSyphilis]. This tree has $75$ leaves, each representing a bacteria strain annotated by: (1) antibiotic resistance ($3$ colors: sensitive, resistant, or n/a), and (2) SNP state ($10$ colors: G, A, T, C, N, each supporting or non-supporting) at a genomic position; there are in total $4$ positions. We applied our algorithm $5$ times; once using the antibiotic resistance coloring, and $4$ using the SNP state colorings. This yielded $5$ (mode, frequency) pairs per clade (i.e., internal node that represents similar bacteria strains indicating transmission between patients). The pairs were used to annotate the tree; a part of the tree annotated only by antibiotic resistance and SNP state for one position is in Fig. 13. By inspecting the (entire) annotated tree, a domain expert (the fourth author of this paper) identified SNPs whose allele patterns correlate with clades of mostly sensitive or mostly resistant strains. This revealed associations worth further statistical testing (e.g., to find out which mutations corresponding to non-supporting alleles cause antibiotic resistance) [10193, Arora2017OriginSyphilis]. Our algorithm took less than a second to produce the annotation information, while manual computation of this information with the aid of a state-of-the-art tool [itol6] took about $2$ hours.
Figure 13: Part of the tree of [10193, Fig. S1] annotated with antibiotic resistance status and SNP state for position $347822$.